My friend stella used to work in a taxi call centre . On any normal day she would expect to take 6calls every 10minutes.

  1. What is the probability that she takes 6calls in the next 10 minutes
  2. she takes a short break between 1023 and 1028 what is the probability that she misses no calls in this period.
1
Expert's answer
2021-07-12T15:04:22-0400

Solution:

Given, "\lambda=6\ calls/10\ min=0.6\ calls/ min"

"X\sim Poi(\lambda)"

(1):

"P(X=6)=e^{-6}\times\dfrac{6^6}{6!}=0.1606"

(2):

Between 10:23 and 10:28, the time duration is 5 minutes.

So, "\lambda=3\ calls/5\ min"

"P(X=0)=e^{-3}\times\dfrac{3^0}{0!}=0.04978"

Now, required probability"=1-0.04978=0.95022"